MOA-II Obsevation Log Monthly Summary 2006/8
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 3093-3942 | 848 | 4944 | 295030 |
LMC | 561-596 | 35 | 355 | 99300 |
SMC | 156-185 | 30 | 211 | 55650 |
ALERT | 326-345 | 20 | 20 | 800 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 933 | 5530 | 450780 (125h13m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
176 | 225 | 174 | 190 | 796 | 163 | 160 | 195 | 795 | 207 | 154 | 153 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
154 | 151 | 147 | 148 | 145 | 144 | 187 | 162 | 159 | 159 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
22 | 19 | 19 | 17 | 17 | 25 | 31 | 32 | 31 | 30 | 25 | 26 |
Field13 | Field14 | Field15 | Field16 |
27 | 28 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
29 | 26 | 24 | 22 | 22 | 22 | 22 | 22 | 22 | 0 | 0 | 0 |
SPECIAL TARGET NAME
Category | Target | Nrun |
ALERT | gb9 | 10 |
ALERT | gb12 | 10 |
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2006/8/1: Alart target MOA 2006 BLG116 (gb10-4) was observed every 10 min. After the second dark, chip 7 is unstable with noise, D1235-D1238 can not be used. We try ./@end and ./setup, after that chip 7 is stable. / night time = 44220(sec)
2006/8/2: Dark D1240 and D1241 can not be uesd due to the (thermal?) noise in chip3 and chip7 . Dark1243 is OK which is taken after ./@end and ./setup. / night time = 44100(sec)
2006/8/3: Alart on the second caustic crossing at OGRE-2006-BLG-277 ( gb10-7 (x,y)=(817.339,2471.795) ). At the DARK, 60sec dark, D1245-60 at the start and D1251-60 at the end is OK.. At the 300sec dark, D1246 and D1247 is OK. D1244 exits the transvers line in chip 6-10. In D1249, chip7 is very bad. One possiblity at the bad DARK is the saturation due to dawn just before fram. / night time = 43980(sec)
2006/8/4: Hifh amplificationnAlart MOA-2006-BLG-126 (fg8, chip10), and OGLE-2006-BLG-437(gb9) is observed every 11 min. Elongation between the field and moon is about 25 degree. DARK D1252-D1253 is OK. / night time = 43860(sec)
2006/8/5: Field gb9 and 12 was observed as Alert A326-345. The exposure time to the bulge is 40 sec. The machine trouble happened, the gears axies of rotator cener breakdown. Then stopped the observation. / night time = 43680(sec)
2006/8/6: The operation is now stopped due to the Machine trouble. The snowfall is a few cm. / night time = 43560(sec)
2006/8/7: Snowfall is 10cm. / night time = 43440(sec)
2006/8/8: The operation will be start soon. Dark D1256-40, D1259-60, D1260-60, D1261-150 are good. / night time = 43320(sec)
2006/8/9: The rotater censer is fixed. The operation restart today! However B3216-B3230 is 3 degree anti-clockwise rotation, and B3231-B3235 is 1.6 degree clockwise rotation due to miss of offset. After B3236 is OK (however same shift renaine. We will reset the offset as soon as possible). 60 sec Dark D1262-1264 and 150sec Dark D1265 are OK. / night time = 43200(sec)
2006/8/10: At 7-8,setting parallel offset value.60sec dark D1266 D1268 are OK and 150sec dark D1269 is OK, / night time = 43080(sec)
2006/8/11: In the dark, chip7 is noisy and unstable sometime. 60 sec dark D1212 and D1273 are OK.150sec dark D1276 is OK. / night time = 42960(sec)
2006/8/12: At the taking dark, chip 7 sometime become a low gain. 60 sec DARK D1278 and 1280 are OK. 15 sec DARK D1283 is OK. In the Flat, F1045-F1059 ,F1061-F1062, F-1064-F1068. and F1070-F1097 are OK. / night time = 42840(sec)
2006/8/13: D1284 , D1286 and D1287 are OK. / night time = 42720(sec)
2006/8/14: 60 sec dark D1289 a,b and D1290 a,b,c are OK. 300sec dark D1291 is ok. Before observation, some troble in cam3 happen, stop at @read. / night time = 42540(sec)
2006/8/15: 60sec dark D1295 and 30 sec dark D1297 300 dark D1296 are OK. MOA 2006 BLG 132 (gb2) and 148 (gb4) are alerted. / night time = 42420(sec)
2006/8/16: 60 sec dark D1298 is OK ,300 sec dark D1299 is OK and CCD tempeature is higher than usual. / night time = 42240(sec)
2006/8/17: 60 sec DARK D1302 and D1308 are OK ,30 sec dark D1303 and D1306 are OK , 15 sec dark D1309 is OK ,300sec dark D1307 is OK. We took flat image. / night time = 42120(sec)
2006/8/18: 60sec dark D1313 is OK, 30sec dark D1311 is OK, 300sec dark D1314 is OK. / night time = 42000(sec)
2006/8/19: 60sec dark D1316 is OK, 30sec dark D1317 is OK. / night time = 41820(sec)
2006/8/20: 60 sec dark D1318 is OK, 300sec dark D1321 is OK / night time = 41700(sec)
2006/8/21: 60s dark D1322 is OK,300sec dark D1323 is OK. / night time = 41520(sec)
2006/8/22: 60sec dark D1326 is OK, 300sec dark D1327 is OK. / night time = 41400(sec)
2006/8/23: 60sec dark D1329 is OK. 300sec dark D1330 is OK. / night time = 41220(sec)
2006/8/24: 60sec dark D1333 is OK. 300sec dark D1334 is OK. / night time = 41100(sec)
2006/8/25: 60sec dark D1336 is OK. / night time = 40920(sec)
2006/8/26: 60sec dark D1340 is OK. 300sec dark D1341 is OK. / night time = 40740(sec)
2006/8/27: 60 sec dark D1342 is OK, 300sec dark D1347 is OK / night time = 40620(sec)
2006/8/28: D1349(60sec), D1350(300sec) are OK. Followed up for MOA151 on gb10 from 1:52am until 2:20am, end of observing for GB. / night time = 40440(sec)
2006/8/29: 60sec dark D13532 is OK,300sec dark D1353 is OK / night time = 40260(sec)
2006/8/30: D1356(60sec) , D1357(300sec) are OK. / night time = 40140(sec)
2006/8/31: D1358(60sec), D1361(300sec) are OK. / night time = 39900(sec)
TOTAL NIGHT TIME: 1308060(sec)
OPERATING RATE: 51.4 %
GB: 37.7 %
LMC: 8.7 %
SMC: 4.9 %
ALERT: 0.1 %
GRB: 0.0 %
Return to index
Previous month
Next month